# OpenVLThinker: An Early Exploration to Complex Vision-Language Reasoning via Iterative Self-Improvement

### v1

@@ 3 Preliminaries
4 OpenVLThinker: Iterative Self-improvement via SFT and RL
We first present the evaluation results of our model OpenVLThinker-7B. Our evaluation employs exact matching and a grader function from MathRuler [^1] https://github.com/hiyouga/MathRuler. We use the same inference hyperparameter as suggested by Qwen and recovered Qwen2.5-VL-7B’s reported results on MathVista. With three iterations of SFT and RL conducted on small amount of data, we observed effective R1-style reasoning and consistent performance gains across several visual reasoning benchmarks including MathVista [33], MathVerse [64] and MathVision [53]. Detailed comparison results are provided in Table 1. As concurrent models remain largely in active development, we exclude these models in our preliminary comparisons. Our evaluation uses the inference hyperparameters as suggested by Qwen and detailed in Table 6.
In this section, we introduce our iterative approach to enhancing complex reasoning capabilities in OpenVLThinker-7B with SFT and RL. Our method consists of three main phases: (1) constructing reasoning-rich datasets through image captioning and distillation from text-based models, (2) using SFT to guide the base model in learning structured reasoning, and (3) applying RL-based exploration to improve generalization. Steps (2) and (3) are conducted iteratively, enabling continuous self-improvement. After each iteration, the RL-fine-tuned model serves as an enhanced source of verified reasoning demonstrations. These improved demonstrations are then used to train the base model through another round of SFT. Additionally, we progressively incorporate more challenging datasets throughout successive iterations, as summarized in Table 2.
Table 1: Evaluation results across multi-modal reasoning benchmarks including MathVista, MathVerse and MathVision. We include the reported performance of GPT-4o as a reference. OpenVLThinker-7B consistently and effectively improves upon the performance of Qwen2.5-VL-7B, surpassing or matching the performance of GPT-4o.
Benchmark GPT-4o (reported) Qwen2.5-VL-7B OpenVLThinker-7B
MathVista 63.8 68.5 70.2
MathVerse 50.2 46.8 47.9
MathVision (testmini) - 27.6 29.6
MathVision (full) 30.4 24.0 25.3
Table 2: Data sources and amounts for different iterations. We progressively evolve the data so that more challenging data are included in later iterations.
Method Data Sources Data Amount
SFT-Iter1 FigureQA, GEOS, Geometry3K, TabMWP, VizWiz, and AI2D 25k
GRPO-Iter1 Geometry3K, GEOS, FigureQA, CLEVR, and SuperCLEVR 5k
SFT-Iter2 FigureQA, GEOS, Geometry3K, TabMWP, VizWiz, AI2D, CLEVR, SuperCLEVR, IconQA, MapQA, and ScienceQA 5k
GRPO-Iter2 Geometry3K, GEOS, FigureQA, CLEVR, and SuperCLEVR 5k
SFT-Iter3 FigureQA, GEOS, Geometry3K, TabMWP, VizWiz, AI2D, CLEVR, SuperCLEVR, IconQA, MapQA, and ScienceQA 5k
GRPO-Iter3 Geometry3K (larger proportion), GEOS, FigureQA, CLEVR, and SuperCLEVR (larger proportion) 5k
4.1 Captioning and Text-based Reasoning Distillation
We curate a collection of visually grounded tasks, including FigureQA [22], GEOS [45], Geometry3K [30], TabMWP [34], VizWiz [3], and AI2D [15], which provide approximately 59.2k image-question-answer triples. These data sources encompass a wide variety of real-world visual tasks. To generate descriptions for each image, we employ a captioning model, Qwen2.5-VL-3B-Instruct, which is prompted to produce textual captions as detailed as possible. Subsequently, a text-based R1 reasoning model, DeepSeek-R1-Distill-14B, is prompted with captions as substitute to images to generate $k=4$ candidate reasoning chains. We select the shortest reasoning path with the correct final answer based on verification with both exact match and the grade function grade_answer() from MathRuler. The generated reasoning length distribution is shown in Figure 10. We verify the model’s final answer with ground-truth answer. We further observe that we obtain fewer reasoning examples with correct answers from challenging tasks like Geometry3K, but more from simpler datasets like TabMWP, as shown in Figure 11. This could be partially due to the information loss during the image-to-text conversion.
4.2 SFT to Follow Reasoning Structure
We initialize our experiments with the base model as Qwen2.5-VL-7B-Instruct, which has broad vision-language pretraining but lacks complex multi-step reasoning abilities. We refer to the initial SFT phase as SFT-Iter1. In this phase, the model is trained on (image, question, reasoning) triplets derived from the data verification procedure and learns to produce the distilled reasoning traces and final answer in a structured format. We monitor performance on the MathVista [33] benchmark to evaluate progress.
We observe that the raw reasoning traces could be excessively verbose or repetitive, likely due to information loss during the image-to-caption conversion process. Consequently, the reasoning behaviors after SFT become longer and even more repetitive, often exhibiting unhelpful self-reflections. An example can be found in Appendix C. To mitigate this issue, we experiment with two filtering strategies. First, we remove samples containing reasoning traces exceeding 500 words. Second, and ultimately adopted, we truncate reflections by identifying and splitting on specific keywords (e.g., “Wait,” “But wait,” “But the question”) and discarding subsequent segments while retaining the final answer, so that we try to prevent the model from learning repeated reflections that significantly degrades its performance. This strategy resulted in 25k shorter examples for training. In Table 3, we compare the performance of model SFTed on our original data against our processed data. We note that although our data cleaning process substantially reduced instances of the "aha moment" behavior from the SFT dataset, such behaviors still occur in our SFT-Iter1 model. Nevertheless, their frequency is considerably more appropriate compared to training with unfiltered SFT data. Below, we present an example of an "aha moment" that persists in the SFT-Iter1 model despite removing all occurrences of the keyword “But wait” in the SFT dataset. [^2] We note that the self-reflective behavior has been observed with text-based Qwen models without distillation as well [63, 29]. The corresponding image for this example is provided in Figure 8.
@@ 4.2 SFT to Follow Reasoning Structure
Table 3: Performance on the MathVista benchmark comparing different SFT data-filtering strategies.
@@
Qwen2.5-VL-7B (baseline) 68.5
Vanilla SFT data (unfiltered) 48.4
Filtered by length ($<$500 words) 55.0
Removed reflections (SFT-Iter1) 62.5
Table 3 shows that unfiltered data severely degraded performance to 48.4%. Restricting the text to under 500 words improved accuracy to 55.0%, while removing repetitive self-reflections further raised it to 62.5%. Consequently, the “Removed reflections” variant served as our dataset for SFT-Iter1. We start from a decent reasoning model and iteratively leverage SFT and RL with a small amount of data to improve model performance.
"But wait" at SFT-Iter1 with correct answer Instruction Move the ruler to measure the length of the line to the nearest centimeter. The line is about (_) centimeters long. SFT-Iter1 <think> Okay, so I need to figure out the length of the line in the image using the ruler provided. Let me start by understanding the image. The image shows a wooden ruler with measurements from 0 cm to 10 cm. The line starts at the 1 cm mark and ends just before the 5 cm mark. Hmm, so the line starts at 1 cm and goes up to just before 5 cm. That means the line is 4 cm long because from 1 cm to 5 cm is 4 cm. But wait, the question is asking for the length to the nearest centimeter. Since the line ends just before 5 cm, it’s not exactly 4 cm, but it’s close. So, rounding to the nearest whole number, it would be 4 cm. </think> The line starts at 1 cm and ends just before 5 cm, which is 4 cm long. Rounding to the nearest centimeter gives 4 cm. <answer>4</answer>
The Role of SFT.
Given the observed performance decline and substantial increase in response length when applying SFT to unfiltered data, we hypothesize that SFT primarily establishes reasoning structures and effectively constrains the search space for RL during subsequent training. Meanwhile, RL training itself should be viewed as the significant contributor to performance gains. While recent studies have explored cold-starting RL training, SFT provides an advantageous warm-start, reducing the need for incorporating format rewards or special prompts that could potentially induce reward hacking. Consequently, we opted for a smaller, carefully processed SFT dataset. Additionally, we report in Table 4 the evaluation results on MathVista for a concurrent work, R1-Onevision [^3] https://huggingface.co/Fancy-MLLM/R1-Onevision-7B [57], which use SFT with approximately 110k visual reasoning data. Under the same evaluation, we similarly observed a performance decline for R1-Onevision.
Table 4: Accuracy (%) on MathVista. While our evaluation recovered the reported performance of Qwen2.5-VL-7B, the performance of SFT-ed model R1-Onevision with 119k SFT data exhibited similar performance decline as our SFT-Iter1 under the same evaluation.
Qwen2.5-VL-7B R1-Onevision-7B SFT-Iter1
MathVista 68.5 61.5 62.5
4.3 Reinforcement Learning with GRPO
After completing SFT-Iter1, we applied GRPO to further improve the model. For this stage, we drew upon a smaller but more challenging subset of 5k examples from Geometry3K, GEOS, FigureQA, CLEVR [21], and SuperCLEVR [25], for which we weren’t able to obtain much correct SFT data via captioning and text-based reasoning. We considered a simple binary reward function (0/1) based on the correctness of the final answer compared to ground truth. No additional template-based or formatting rewards were used, as SFT had already enabled the model with a structured reasoning style.
During training, in Figure 5, we observed a correlation between longer reasoning outputs and improved accuracy on a held-out test set, while verbose self-reflections were markedly less frequent than the SFT model trained with unfiltered data. On MathVista, GRPO-Iter1 improved the accuracy from 62.5% (SFT-Iter1) to 65.6%, suggesting that RL can effectively improve the model’s reasoning capability once the basic reasoning structure is established.
(a) Average response length
(b) Test score
Figure 5: Training progress with GRPO at iteration 1. We report the test accuracy on a held-out validation set from the same distribution as the training data.
4.4 Iterative Refinements
The model obtained from our first iteration (GRPO-Iter1) now serves as a better source of reasoning demonstrations, having acquired complex reasoning capabilities while more reliable on processing visual inputs as compared to image-text conversion. This motivates us to adopt an iterative self-improvement strategy, which is proven to be effective in iterative SFT [48] and iterative DPO schemes [60, 41]. Following the initial RL phase, we generated a new set of higher-quality reasoning traces by sampling from the GRPO-Iter1 model. This yielded another 5k examples drawn from more diverse sources, including FigureQA, GEOS, Geometry3K, TabMWP, VizWiz, AI2D, CLEVR, SuperCLEVR, IconQA [31], MapQA [4], and ScienceQA [32]. We referred to the subsequent SFT round on these data as SFT-Iter2. Empirically, SFT-Iter2 with less and higher-quality data (5k) further reduced the base model’s performance decline on MathVista, reaching 66.1%.
We then repeated the GRPO process on SFT-Iter2, denoted GRPO-Iter2, using the same 5k curated examples from the previous RL stage. Notably in Figure 6, although the average response length decreased in this second RL round, the accuracy on MathVista increased to 69.4%, surpassing the original baseline of 68.5%. This confirms that iterative cycles of supervised fine-tuning and reward-based optimization can progressively enhance performance. We summarize the model’s performance on MathVista across iterations in Figure 4.
(a) Average response length
(b) Test score
Figure 6: Training progress with GRPO at iteration 2. We report the test accuracy on a held-out validation set from the same distribution as the training data.

### v2

@@ 3 Preliminaries
4 OpenVLThinker: Iterative Self-improvement on Curriculum Data
Figure 5: Curation of SFT-Iter1 data from text-based reasoning models based on image descriptions.
In this section, we first analyze how SFT and RL affect the occurrence of reasoning-related keywords, which serves to motivate our approach. We then introduce the proposed iterative approach to enhancing complex reasoning capabilities in OpenVLThinker-7B with SFT-RL cycles. At last, we propose a source-based curriculum RL.
4.1 The Role of SFT and RL
The initial SFT data.
The standard distillation approach used for text-only reasoning cannot be directly applied because the R1 model does not support visual input, and other proprietary LVLMs, such as OpenAI’s o1/o3, do not expose their internal reasoning paths. To learn reasoning behaviors from R1, we instead use the target model as a captioning model, prompting it to generate detailed textual descriptions for each image. Subsequently, these captions serve as proxies for the images when input into a text-based R1 reasoning model, QwQ-32B [61], which then generates $k$ candidate reasoning chains. Among these candidates, we select the shortest reasoning chain that correctly arrives at the final answer to avoid excessive reasoning length after SFT (further details in Section 5.2). The overall procedure is summarized in Figure 5.
Impact of SFT and RL on Model Reasoning Actions.
Complex reasoning behaviors in LLMs have been described using various terms, including long CoT [77] and aha moments [19]. At their core, these behaviors reflect autonomous planning, reflection, and verification steps that occur during inference. We refer to them as inference-time actions, which are often triggered by specific tokens such as “wait”. To examine how SFT and RL influence these reasoning actions, we identify eight representative keywords corresponding to perception, question comprehension, planning, reflection, and seeking alternatives.
Figure 6: Occurrences of reasoning keywords when solving MathVista with the base model, SFT-Iter1 model, and GRPO-Iter1 model. The most significant distribution shift occurs after SFT, while the scale remains largely unchanged after GRPO, despite notable performance improvements.
As illustrated in Figure 6, the base model seldom exhibits planning, reflection, or alternative-solution actions. However, SFT guided by text-based R1-like reasoning traces effectively surfaces these behaviors. As shown in the third and fourth subplots of Figure 6, subsequent GRPO-based RL training following SFT-Iter1 substantially enhances model performance on MathVista by 5.2%, yet largely maintains the initial reasoning action distribution, with minor refinements such as reduced repetitive reflections.
Conversely, direct RL training without prior SFT struggles to efficiently induce reasoning behaviors, exemplified by the absence of reflection keywords (e.g., “wait”) even after an equivalent training volume. Concurrent research by [66], which solely relies on RL, addresses this by explicitly appending relevant keywords during training rollouts. These observations support our argument that SFT plays a critical role in highlighting desirable reasoning actions, providing an efficient and effective foundation for RL to build upon. In contrast, RL primarily serves to further refine and enhance performance.
4.2 Iterative Improvement
The model obtained after the first iteration (GRPO-Iter1) demonstrates enhanced complex reasoning capabilities and improved reliability in processing visual inputs compared to methods based on image-to-text conversion. This advancement positions GRPO-Iter1 as an effective source for generating higher-quality reasoning demonstrations. Consequently, we propose an iterative self-improvement strategy, inspired by established methodologies such as iterative SFT in ReST-EM [57] and iterative direct preference optimization (DPO) schemes [79, 50], both of which have shown substantial effectiveness in iterative training processes and fall under the Expectation-Maximization framework [57].
Specifically, in each iteration, we sample a new set of enhanced reasoning traces using the model trained in the preceding iteration. These refined demonstrations are then utilized to retrain the base model [^1] To maintain stability, we retrain the model from scratch at each iteration with the newly generated dataset, as similar to some iterative approaches in text-only domain [57, 23]., thereby progressively elevating its reasoning performance. The overall iterative pipeline is illustrated in Figure 2, and the consistent incremental performance gains achieved through successive iterations are depicted in Figure 4.
4.3 Two-Stage Source-Based Curriculum RL
To ensure effective exploration during reinforcement learning (RL), we assess the difficulty of data sources, aiming to provide data that is challenging yet appropriate for the model’s proficiency level. Specifically, we utilize GPT-4o to rate the difficulty of five representative examples drawn from various data sources such as FigureQA [30], MapQA [4], and GeoQA [7], in a similar fashion to the text-based evaluation in DeepMath-103K [22]. Additionally, we employ the base model, Qwen2.5-VL-7B [^2] In alignment with previous R1 reasoning research [77, 76], we choose the base model from Qwen2.5 family for their strong general capability obtained in pre-training., to obtain its error rates as a complementary difficulty indicator. We standardize independently using z-score normalization for both the GPT-4o rating and base model error rates and compute the average of the two. Based on this composite score, we categorize the data sources into Easy, Medium, and Hard groups via k-means clustering in 1d space. With these categories, we construct two difficulty-specific datasets: $\mathcal{D}_{\mathrm{RL(Medium)}}$ and $\mathcal{D}_{\mathrm{RL(Hard)}}$. Our curriculum training thus proceeds in two stages within one iteration, sequentially training on $\mathcal{D}_{\mathrm{RL(Medium)}}$ and $\mathcal{D}_{\mathrm{RL(Hard)}}$.
5 Experiments
Training setup. We take Qwen2.5-VL-7B [3] as the base model and perform three iterations of the SFT-RL cycle as illustrated in Section 4, applying full fine-tuning for both SFT and RL. Our training framework is based on LLaMA-Factory [^3] https://github.com/hiyouga/LLaMA-Factory for SFT and EasyR1444https://github.com/hiyouga/EasyR1 for RL. We source our training data from the established LLaVA-OneVision [33] and specifically consider the 14 data sources in overlap with MathV360K [56] (Table 4). Based on our preliminary experiments, we equally draw 500 examples from each source to form the SFT seed dataset of 7K examples, where for each iteration we collect distillation data via rejection sampling, resulting in a final 3K SFT data. We then classify the data sources into easy, medium and hard (as detailed in Table 4). We construct the 3K medium-level RL training data from the 5 sources that we identified as medium difficulty. Finally, we construct 6K hard-level RL training data from the 3 most difficult sources, summing up to 12K data in total for each iteration that trains from the base model. We defer the training hyperparameters to Appendix B.
Evaluation. Our evaluation employs exact matching and a grader function from MathRuler [^5] https://github.com/hiyouga/MathRuler. We use the same inference hyperparameter as suggested by Qwen and recovered Qwen2.5-VL-7B’s reported results on MathVista at 68.5%. The hyperparameters are detailed in Table 6. We employ six established benchmarks to examine model’s ability thoroughly:
  • Math reasoning: MathVista [43], MathVerse [86] and MathVision [67]. The three benchmarks evaluate how LVLMs interpret and reason with diagrams in visual math problems through both multiple-choice and free-form questions.
  • General reasoning: MMMU-Pro [80] and EMMA [21]. MMMU-Pro spans 30 subjects across 183 subfields, including business, medicine, and science. EMMA evaluates in physics, chemistry, coding, and math.
  • Perception: HallusionBench [18], designed to evaluate LVLMs’ susceptibility to language hallucination and visual illusion.
Baselines. We evaluate the non-reasoning base model Qwen2.5-VL-7B as a primary baseline to demonstrate the improvements introduced by our method. Additionally, we include the reported performance of proprietary models, including GPT-4o [28] and Claude-3.5-Sonnet [2], alongside open-source LVLMs such as Mulberry-7B [75], InternVL2.5-8B [8], Kimi-VL-16B [59], and Qwen2-VL-7B [68], as reference points. Crucially, to highlight the effectiveness of our iterative SFT-RL training strategy, we compare our model with concurrent approaches employing a single round of SFT distillation and RL at the same model scale (7B), yet utilizing significantly larger training datasets. These concurrent models include R1-VL-7B [84], R1-Onevision-7B [74], and VLAA-Thinker-Qwen2.5VL-7B [6]. Notably, R1-Onevision and VLAA-Thinker-Qwen2.5VL-7B also start from the same base model (Qwen2.5-VL-7B) as ours, using 165K and 150K total data, respectively. In contrast, our model achieves better performance with only 12K training samples from the base model.
5.1 Main Results
Table 1: Evaluation results across visual math reasoning benchmarks (MathVista, MathVerse, MathVision), general visual reasoning benchmarks (MMMU-Pro, EMMA), and perception (HallusionBench). We include the reported performance of proprietary models and open-source Vision-Language models as references. *Performance of the base model Qwen2.5-VL-7B and concurrent reasoning models are evaluated by us under the same setting and hardware as OpenVLThinker. The bold numbers indicate the best results among the open-source models and the underscored numbers represent the second-best results.
Math Reasoning General Reasoning Visual
Model Data Math-Vista Math-Verse Math-Vision MMMU-Pro EMMA Hallu-Bench Avg
Proprietary Model
GPT-4o - 63.8 50.2 30.4 51.9 32.7 55.0 47.3
Claude-3.5-Sonnet - 67.7 47.8 33.5 51.5 35.1 55.0 48.4
Open-source Vision-Language Model
Mulberry-7B - 63.1 39.6 - - - 54.1 -
InternVL2.5-8B - 64.4 39.5 19.7 34.3 - - -
Kimi-VL-16B - 68.7 44.9 21.4 - - - -
Qwen2-VL-7B - 58.2 32.5 16.3 30.5 20.2 50.6 34.7
Qwen2.5-VL-7B* - 68.5 46.8 24.0 42.3 24.4 67.5 45.6
Concurrent Vision-Language Reasoning Models
R1-VL-7B 270K 63.5 40.0 24.7 7.8 8.3 54.7 33.2
R1-Onevision-7B 165K 64.1 46.4 29.9 21.6 20.8 65.6 41.4
VLAA-Thinker-7B* 150K 70.0 48.6 24.9 42.0 25.5 68.4 46.6
OpenVLThinker-7B* 12K 72.3 50.3 25.9 42.9 26.8 69.1 47.9
Table 2: Performance of 3B models on MathVista.
Model Accuracy (%)
R1-VL-2B 52.1
InternVL2.5-4B 60.5
Qwen2.5-VL-3B 62.3
VLAA-Thinker-3B 61.0
OpenVLThinker-3B 63.4
We present our main results in Figure 1, with detailed performance across datasets shown in Table 1. As illustrated, OpenVLThinker-7B consistently achieves either the best or second-best scores among open-source LVLMs of comparable scale across all six benchmarks, including concurrent reasoning models. On average, OpenVLThinker attains an accuracy of 46.6%, representing a 2% improvement over the base model and performance comparable to proprietary models such as GPT-4o. Notably, OpenVLThinker exhibits fewer hallucinations and more precise perception than its base model on HallusionBench, improving accuracy by 2.7%. Compared to concurrent reasoning methods that utilize substantially larger datasets for single-iteration SFT and RL, our iterative approach achieves superior results while utilizing only 1/10 of the data scale as used in concurrent works with a single-iteration SFT-RL pipeline.
OpenVLThinker-3B. We additionally train a 3B model using a single iteration of the SFT-RL pipeline, where the training process distills from our 7B model. In Table 2, we compare the performance of our 3B model against current representative models at the same scale, including our base model, Qwen2.5-VL-3B, and the reasoning model VLAA-Thinker-3B, which is trained from the same initial checkpoint as ours. OpenVLThinker-3B achieves the best performance on MathVista and outperforms state-of-the-art 3B reasoning models.
5.2 Analysis
Distillation at iteration 1.
At SFT-Iter1, we utilized the base model Qwen2.5-VL-7B to generate image descriptions and obtained R1-like reasoning from QwQ-32B through rejection sampling. A common problem for distillation observed in text-only math reasoning is the overly long reasoning length coupled with unnecessary repetitions of reflections [77, 41]. We observed similarly that these initial reasoning traces were often excessively verbose, partly due to information loss during image-to-caption conversion. Consequently, post-SFT reasoning became increasingly repetitive with unproductive self-reflections (see Appendix C for an illustration). To address this, we evaluated two filtering strategies: (1) discarding samples with reasoning traces exceeding 500 words, and (2) truncating reflections by splitting traces of at specific keywords that were overly repetitive in data ( “Wait,” “But wait,” and “But the question”) and discarding subsequent segments while preserving the final answer. The latter approach was ultimately adopted to prevent the model from internalizing reflection loops, while preserving the reasoning action at a reasonable frequency. Table 5.2 compares models trained on original versus processed data.
@@ 5.2 Analysis
Table 3: Performance on the MathVista benchmark comparing different SFT data-filtering strategies. Removing the most repetitive keywords in data can mitigate repetitive reflections after SFT.
@@
Qwen2.5-VL-7B 68.5
Vanilla 57.5
Filtered 58.7
Truncated (SFT-Iter1) 63.4
Table 4: Categorization of data sources by composite difficulty score using k-means with k=3. The geometry question sources all fall into the hard category.
Easy Medium Hard
ChartQA FigureQA UniGeo
IconQA CLEVR GEOS
VizWiz A-OKVQA Geometry3K
TabMWP SuperCLEVR GeoQA
DVQA MapQA
Data source difficulty.
We conducted a quantitative analysis to categorize the data sources based on difficulty. Applying k-means clustering (with $k=3$) to our composite difficulty score as described in Section 4 allowed us to clearly identify three distinct difficulty pools, as shown in Table 4. We visualize the difficulty scores for each source in Figure 7. In Figure 8, we show the performance of GRPO-Iter1 when drawing 3K data from either (1) 10 data sources classified as either Easy or Medium, or (2) 5 data sources classified as Medium. We observe that RL training with easy-level data results in ineffective performance gain as compared to sourcing from medium-level data only. This finding aligns with concurrent algorithmic efforts such as DAPO [78] in the text-only domain for improving GRPO by dynamically filtering out overly-easy examples.
Figure 7: Data source difficulty based on base model accuracy and GPT-4o rating.
Figure 8: Performance at GRPO-Iter1 using data from different difficulty sources, at the same scale of 3K.
Curriculum RL to maximize utilization of challenging data.
Figure 9 investigates the impact of incorporating challenging training data (e.g., geometry datasets) at iteration 1. On the left panel, we illustrate the absolute performance gains transitioning from SFT-Iter1 to GRPO-Iter1 (medium difficulty), and subsequently from GRPO-Iter1 (medium) to GRPO-Iter1 (hard). Training on these harder datasets yields substantial improvements on more difficult benchmarks, such as MathVision, while not significantly affecting performance on easier benchmarks like MathVista. On the right panel, we further compare our two-stage, source-based curriculum RL approach against training solely on hard data. The results indicate that initiating RL with moderately challenging (medium difficulty) data and subsequently progressing to harder datasets provides optimal performance improvements.
Figure 9: Absolute performance gain achieved at iteration 1. The round 2 RL training on hard data provides more significant performance gain on harder benchmarks such as MathVision. Moreover, if RL training with the hard data only yield less improvement than our two-stage curriculum RL.
Figure 10: Distribution of reasoning length (number of words) across iterations of training. While our trained reasoning model across iterations all tend to reason longer than the base model, iterative training resulted in gradually more concise length, possibly due to reduced repetitive reflections.
Iterative progression.
Building upon the performance improvements shown in Figure 4, we further analyze changes in reasoning length across iterations, as illustrated in Figure 10. Our results indicate that the reasoning model consistently utilizes more words at inference time compared to the base non-reasoning model, without becoming excessively repetitive. Notably, the largest increase in reasoning length occurs at Iteration 1, with subsequent iterations gradually adopting more concise reasoning. This progression suggests an increasingly efficient utilization of reflective reasoning, engaging reflections primarily when beneficial. In Appendix C (Figure 13 and 14), we show reasoning examples that our SFT-ed model was incorrect while our RL-ed model was correct.